40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 24 18:21:24 2011, kiwamu, Update, Green Locking, in-loop and out-of-loop measurements green_lock_setup.pngmeasurement1.pngTRX_REFL11.pngmeasurement2.png
    Reply  Thu Feb 24 19:59:25 2011, kiwamu, Update, Green Locking, whitening filter for ALS ACDC.png
Message ID: 4352     Entry time: Thu Feb 24 18:21:24 2011     Reply to this: 4353
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: in-loop and out-of-loop measurements 

Two different measurement have been performed for a test of the green locking last night.

Everything is getting better. yes. yes.

green_lock_setup.png

 


[ measurement 1 : IR locking]

The X arm was locked by using the IR PDH signal as usual (#4239#4268) .

The in-loop signal at from the IR path and the out-of-loop signal at from the green beat note path were measured.

measurement1.png

Let us look at the purple curve. This is an out-of-loop measurement by looking at the green beat note fluctuation.

The rms down to 0.1 Hz used to be something like 60 kHz (see here), but now it went down to approximately 2 kHz. Good.

This rms corresponds to displacement of about 260 pm of the X arm. This is barely within the line width. The line width is about 1 nm.

 

 


[ measurement 2 : green locking]

The motion of the X arm was suppressed by using the green beat signal and feeding it back to ETMX.

After engaging the ALS servo, I brought the cavity length to the resonance by changing the feedback offset from epics.

Then took the spectra of the in-loop signal at the beat path and the out-of-loop signal at the IR PDH path.

 

 Here is a time series of TRX after I brought it to the resonance.

TRX_REFL11.png

TRX was hovering around at the maximum power, which is 144 counts.

measurement2.png

Since I put one more 10:1 filter to suppress the noise around 3 Hz, the rms of the in-loop beat spectrum went to about 1 kHz, which used to be 2 kHz (see #4341).

But the out-of-loop (IR PDH signal) showed bigger noise by a factor of 2 approximately over frequency range of from 2 Hz to 2 Hz. The resultant rms is 2.7 kHz.

The rms is primarily dominated by a peak at 22 Hz (roll mode ?).

I calibrated the IR PDH signal by taking the peak to peak signal assuming the finesse of the cavity is 450 for IR. May need a cooler calibration.

ELOG V3.1.3-